The quantity of goods and prices in each year are given

Chicken tomato Park Ticket
2001 Price Rs.80 Rs.24 Rs. 5.00
2001 Quantity 40 1 4


Now, 20 years later, we have new prices:
2021 Chicken tomato Park Ticket
Rs 120 Rs 40 Rs 12

Find out the inflation rate from 2001 to 2021 with the (hint CPI measurement)
1
Expert's answer
2021-04-19T18:51:23-0400

2001:


"80\times40+24\times1+5\times4=3244"

2021:


"120\times40+40\times1+12\times4=4888"

"CPI=\frac{4888}{3244}=1.507"